OpenStudy (linstead_cpd_freak):

What is the solution to the equation? 5x + 10 = –4x – 17 A. -3 B. 3 C. 7 D. -7/9

rebeccaxhawaii (rebeccaxhawaii):

Get like terms on each side. - x's on one and constants on the other.

rebeccaxhawaii (rebeccaxhawaii):

so 5x+4x=-17-10

OpenStudy (linstead_cpd_freak):

Ok

OpenStudy (linstead_cpd_freak):

9x = -27 ? Sorry I'm confused

OpenStudy (linstead_cpd_freak):

@Mehek14 Help please

Mehek (mehek14):

yes 9x = -27 now we isolate x so -27/9 = ? :)

OpenStudy (linstead_cpd_freak):

-3

Mehek (mehek14):

so x = -3

OpenStudy (linstead_cpd_freak):

Thank you

Mehek (mehek14):

You're welcome :)